OpenStudy (anonymous):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2

OpenStudy (anonymous):

2x + y = 2

OpenStudy (anonymous):

(y+2) / (x-2) = -1/2 2y+4 = -x+2 2y+x = -2

OpenStudy (anonymous):

I don't know how @Raya_Atour got her answer?

OpenStudy (anonymous):

4 + -2 = 2 -10 + 12 = 2

OpenStudy (anonymous):

I used the first point and got my answer 2y+x = -2